OpenStudy (anonymous):

The odds against an event are 7:2. Find the probability that the event will occur.

hartnn (hartnn):

odds against an event = (1-p)/p = 7/2 find 'p' from here.

OpenStudy (anonymous):

see if odds against event is m:n then the probability that the event will occur is given by n/(m+n)

OpenStudy (anonymous):

Isnt it p/(1-p) =7/2

OpenStudy (anonymous):

thus here the probability that the event will occur is 2/(2+7)= 2/9

hartnn (hartnn):

p/(1-p) is odds in favor.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

Odds notation always has me running to Google.